OpenStudy (anonymous):

i need help on equations that are equivalent to 0 the problem I'm stuck on is 3x(x+6)=0

OpenStudy (anonymous):

Use the distributive property. 3x*x+ 3x*6=0

OpenStudy (anonymous):

^ That's redundant if you're trying to find what x =

OpenStudy (anonymous):

You would get 3x^2 + 18x=0 Subtract 18x and you get 3x^2=-18x Divide out the x and you get 3x=-18 Divide out the 3 and you get x=-6

OpenStudy (anonymous):

That's completely unnecessary and only half right. You already have your equation in factored form, just equate x to 0